# A structure-preserving semi-implicit four-split scheme for continuum mechanics

### v1

@@ 1 Introduction
The primal objective of this work is to design a novel four-split scheme for the GPR model of continuum mechanics [30] that accounts for both fluid and solid mechanics within one and the same set of equations. It covers a very wide range of phenomena, spanning from ideal and viscous heat conducting fluids to elastic and elasto-plastic solids. Indeed, this is possible thanks to the presence of stiff relaxation sources in the evolution equation of the deformation gradient and of the thermal impulse, which retrieve the Navier-Stokes-Fourier model in the stiff relaxation limit [31]. For vanishing source terms, the mechanical behavior of ideal elastic solids is reproduced, involving both shear and thermal stresses.
@@ 2 Governing PDE system and splitting
The unified first order hyperbolic GPR model of continuum mechanics including heat conduction, see also [67, 30, 31, 68], reads as follows
@@
The system is thermodynamically compatible and satisfies the following entropy inequality:
@@
Throughout this paper we assume that the temperature is positive, i.e. $T>0$. The state vector is given by $\mathbf{q}=\{q_{i}\}=(\rho,\rho v_{i},\mathcal{E},A_{ik},J_{k})^{T}$, with the mass density $\rho$, the velocity field $v_{i}$, the distorsion field $A_{ik}$, the specific thermal impulse $J_{k}$ and the total energy density $\mathcal{E}=\rho E=\mathcal{E}_{1}+\mathcal{E}_{2},+\mathcal{E}_{3}+\mathcal{E}_{4}$ containing four different contributions, namely internal energy, kinetic energy, energy due to the elastic deformation of the medium and a contribution due to the presence of the specific thermal impulse,
@@
In the numerical scheme, we will use also the governing equation for the metric tensor, instead of solving for the distortion field directly. It can be derived from (1c) and reads
@@ 2.3 G-J-v subsystem
The next subsystem concerns the coupling between momentum equation via the two stress tensors with the distortion matrix and specific thermal impulse. Here, we have replaced (1c) by (11) due to the favorable structure of the metric tensor equation. Due to the appearance of the stress tensors in the total energy equation, it is added to this subsystem while the other variables have been omitted.
@@ 3.2 Discrete temperature subsystem
$\mathcal{E}^{**,c}$ $=\mathcal{E}^{*,c}-\partial_{k}^{cp}q_{k}^{**,p}.$ (30b)
@@ 3.3 Discrete $\mathring{G}$-J-v subsystem
In order to obtain $\mathring{G}$-J-v we start from a discrete G-J-v subsystem:
@@
where we used the symmetry of $G$ to simplify the trace contribution. We now introduce the following semi-implicit approximations for the two stress tensors
@@ 3.3.2 Asymptotic limit: The Fourier heat conduction
Moreover, in the discrete G-J-v subsystem by (41) using the Hilbert expansion of $J_{k}^{***,c}$ one obtains immediately $J_{k}^{***,c}=\mathcal{O}(\tau_{2})$. In the pressure subsystem presented in the sequel, the final update $J_{k}^{n+1,c}$ is well-prepared with the analogue argumentation. Therefore, in the stiff relaxation limit $\tau_{2}\to 0$ we recover the discrete Fourier heat flux independently of $\Delta t$.
@@ 3.4 Discrete pressure subsystem
The coefficient matrix of this system is symmetric and positive definite, and (55) can therefore be solved easily via a classical matrix-free conjugate gradient method. The final momentum is then given by (54a) and the scheme is completed by formally setting $\rho^{n+1,p}=\rho^{*,p}$, while the distorsion field and the thermal impulse are updated via the following compatible structure-preserving discretization that preserves curl-free fields exactly at the discrete level in the absence of source terms, by virtue of the use of compatible discrete nabla operators:
@@
In order to enforce the compatibility with the determinant constraint we may uniformly rescale all components of $A_{ik}^{n+1,c}$ so that $|A_{ik}^{n+1,c}|=\rho^{n+1,c}/\rho_{0}$ holds. For alternative compatible discretizations, see [75, 74].
$J_{k}^{n+1,c}=J_{k}^{n,c}-\Delta t\,\partial_{k}^{cp}\left(v_{m}^{n+1,p}J_{m}^{n,p}\right)-\frac{\Delta t}{4}\sum\limits_{p\in\mathcal{P}_{c}}v_{m}^{n+1,p}\left(\partial_{m}^{pa}J_{k}^{n,a}-\partial_{k}^{pa}J_{m}^{n,a}\right)+\frac{\Delta t}{\theta_{2}}\beta_{ik}^{n+1,c}.$ (57)
The discrete total energy is finally updated as
@@ 3.5 Summary of the algorithm
In the following we summarize the complete algorithm of the proposed structure-preserving four-split scheme:
@@ 3.5.2 Temperature subsystem
$\mathcal{E}^{**,c}=\mathcal{E}^{*,c}-\partial_{k}^{cp}q_{k}^{**,p}.$ (63)
@@ 3.5.5 Compatible update of the distortion field, of the thermal impulse and of the total energy
In the absence of source terms the distortion field and the thermal impulse must remain curl-free for all times if they were initially curl-free. This is achieved via
@@
$J_{k}^{n+1,c}=J_{k}^{n,c}-\Delta t\,\partial_{k}^{cp}\left(v_{m}^{n+1,p}J_{m}^{n,p}\right)-\frac{\Delta t}{4}\sum\limits_{p\in\mathcal{P}_{c}}v_{m}^{n+1,p}\left(\partial_{m}^{pa}J_{k}^{n,a}-\partial_{k}^{pa}J_{m}^{n,a}\right)+\frac{\Delta t}{\theta_{2}}\beta_{ik}^{n+1,c}.$ (70)
@@ 4.4 Lid-driven cavity
Refer to caption Refer to caption
@@ 4.5 Solid rotor
The aim of this test case is to numerically verify the discrete curl-free property of our new 4-split semi-implicit finite volume scheme. For the setup of the test case we follow [32], solving (1) with $\tau_{1}=10^{20}$, but here we set $\tau_{2}=10^{-14}$, i.e. the strain relaxation source term is absent, while heat conduction is considered in the stiff Fourier limit. The computational domain is given by $\Omega=[-1,+1]^{2}$. The initial condition is $\rho=1$, $p=10^{5}$, $\mathbf{A}=\mathbf{I}$ and $\mathbf{J}=0$ and the initial velocity field is u=y/Ru=-y/R, $v=+x/R$ and $w=0$ within the circular region $r\leq R$, where $r=\left\|\mathbf{x}\right\|_{2}$ and $R=0.2$, while $\mathbf{v}=0$ for $r>R$. The remaining parameters are $\gamma=1.4$, $c_{v}=1004/\gamma$, $c_{s}=1.0$ and $c_{h}=100$. With this setting we are in the low Mach regime for the acoustic Mach number $M_{a}$ and for the heat Mach number $M_{h}$. The test case is solved with the new structure-preserving semi-implicit 4-split finite volume scheme until a final time of $t=0.3$. The computational mesh uses $512\times 512$ cells. In order to obtain a numerical reference solution, the same problem is solved on the same grid with the second order version of the 2-split structure-preserving scheme introduced in [32], but in the absence of heat conduction, i.e. setting $c_{h}=0$, since the 2-split method presented in [32] is not able to deal with low heat Mach numbers $M_{h}$. The results obtained with the new 4-split scheme are depicted in Fig. 9, where the contour colors of the velocity component $u$ and the distortion field components $A_{11}$ and $A_{12}$ are shown. A comparison with the reference solution is provided in the 1D cut shown in the left panel of Fig. 10, while in the right panel of Fig. 10 we show the the time series of the curl errors of $\mathbf{A}$ and $\mathbf{J}$ in the $L^{\infty}$ norm. One can observe a good agreement with the reference solution and that the new structure-preserving semi-implicit 4-split scheme presented in this paper produces curl errors that remain at the order of machine precision, as expected.
@@ 4.6 Circular explosion problem
We first solve the governing PDE system in the quasi inviscid fluid limit of the model, i.e. $\tau_{1}\ll 1$ and $\tau_{2}\ll 1$. For the inner state we set the density and the pressure to $\rho_{in}=1$ and $p_{in}=1$, while in the outer state we impose $\rho_{out}=0.125$ and $p_{out}=0.1$. The initial velocity is globally set to $\mathbf{v}=0$, the initial thermal impulse vector is equal to $\mathbf{J}=0$ and the distortion field is initialized as $\mathbf{A}=\mathbf{I}$. The other parameters are chosen as $\gamma=1.4$, $c_{v}=2.5$, $c_{s}=1$, $c_{h}=1$, $\rho_{0}=1$, $\tau_{1}=10^{-8}$ and $\tau_{2}=10^{-10}$. The mesh is composed by $500\times 500$ control volumes and the final time is set to $t=0.2$. Note that with the chosen parameters, all Mach numbers are of the order of unity. The reference solution is obtained by solving a 1D radial Euler system with source terms using a classical second-order MUSCL-Hancock scheme on a very fine mesh, as described in detail in [80, 32]. In the left panel of Fig. 11 the 3D density contours are shown at the final time. In Fig. 12 a comparison of the reference solution with the numerical results obtained at the aid of the new semi-implicit 4-split scheme is provided. One observes overall a good agreement.
@@
We now solve the governing PDE system in the solid limit, i.e. $\tau_{1}\to\infty$ and $\tau_{2}\to\infty$, with the same initial conditions and parameters as described above for EP1, apart from the relaxation times, which are set to $\tau_{1}=\tau_{2}=10^{20}$. The mesh is again composed of $500\times 500$ cells and the final time is now set to $t=0.15$. We solve the problem twice, once with the new 4-split SPSIFV scheme and another time with the thermodynamically compatible HTC scheme presented in [68, 82], in order to obtain a numerical reference solution. A 3D density contour plot at the final time is depicted in the right panel of Fig. 11. The comparison of the computational results between the new 4-split SIFV scheme and the numerical reference solution is shown in via 1D cuts along the $x$-axis in Fig. 13. One can again note a good agreement between the two solutions.
@@ 5 Conclusions
Further research will be devoted to the adoption of unstructured meshes with the aim of tackling more complex geometries, while preserving additional structure of the system exactly, see e.g. [75, 73, 74], as well as the incorporation of electromagnetic fields, following the general lines of [83]. To improve the accuracy in space, high order extension is also foreseen following the ideas recently forwarded in [36], while higher order in time may be achieved according to the lines presented in [84].

### v2

@@ 1 Introduction
The primal objective of this work is to design a novel four-split scheme for the Godunov-Peshkov-Romenski (GPR) model of continuum mechanics [30] that accounts for both fluid and solid mechanics within one and the same set of equations. It covers a very wide range of phenomena, spanning from ideal and viscous heat conducting fluids to elastic and elasto-plastic solids. Indeed, this is possible thanks to the presence of stiff relaxation sources in the evolution equation of the deformation gradient and of the thermal impulse, which retrieve the Navier-Stokes-Fourier model in the stiff relaxation limit [31]. For vanishing source terms, the mechanical behavior of ideal elastic solids is reproduced, involving both shear and thermal stresses.
@@ 2 Governing PDE system and splitting
The unified first order hyperbolic model of continuum mechanics of Godunov, Peshkov and Romenski, [67, 30, 31, 68], hereafter also simply denoted by GPR model, reads as follows:
@@
Throughout this paper we make use of the Einstein summation convection over two repeated indices. The system is thermodynamically compatible and satisfies the following entropy inequality:
@@
Throughout this paper we assume that the temperature is positive, i.e. $T>0$. The state vector is given by $\mathbf{q}=(\rho,\rho v_{i},\mathcal{E},A_{ik},J_{k})^{T}$, with the mass density $\rho$, the velocity field $v_{i}$, the distorsion field $A_{ik}$, the specific thermal impulse $J_{k}$ and the total energy density $\mathcal{E}=\rho E=\mathcal{E}_{1}+\mathcal{E}_{2},+\mathcal{E}_{3}+\mathcal{E}_{4}$ containing four different contributions, namely internal energy, kinetic energy, energy due to the elastic deformation of the medium and a contribution due to the presence of the specific thermal impulse,
@@
In the numerical scheme, we will also make use of the governing equation for the metric tensor as auxiliary quantity, since it will allow us to obtain a linearly implicit scheme in the strain relaxation source term that is asymptotically consistent with the Navier-Stokes equations in the stiff relaxation limit when $\tau_{1}\to 0$. It can be derived from (1c) and reads
@@ 2.3 G-J-v subsystem
The next subsystem concerns the coupling between momentum equation via the two stress tensors with the distortion matrix and specific thermal impulse. Here, we have replaced (1c) by (11). Due to the appearance of the stress tensors in the total energy equation, it is added to this subsystem while the other variables have been omitted:
@@ 3.2 Discrete temperature subsystem
$\mathcal{E}^{**,c}$ $=\mathcal{E}^{*,c}-\Delta t\,\partial_{k}^{cp}q_{k}^{**,p}.$ (30b)
@@ 3.3 Discrete $\mathring{G}$-J-v subsystem
In order to obtain the $\mathring{G}-J-v$ system which is actually used in our scheme, we start from a discrete GJvG-J-v subsystem:
@@
where we used the symmetry of $\mathbf{G}$ to simplify the trace contribution. We now introduce the following semi-implicit approximations for the two stress tensors
@@ 3.3.2 Asymptotic limit: The Fourier heat conduction
Moreover, in the discrete G-J-v subsystem by (41) using the Hilbert expansion of $J_{k}^{***,c}$ one obtains immediately $J_{k}^{***,c}=\mathcal{O}(\tau_{2})$. In the pressure subsystems presented in the sequel, the final update $J_{k}^{n+1,c}$ is well-prepared with the analogue argumentation. Therefore, in the stiff relaxation limit $\tau_{2}\to 0$ we recover the discrete Fourier heat flux independently of $\Delta t$.
@@ 3.4 Discrete pressure subsystem
The coefficient matrix of this system is symmetric and positive definite, and (55) can therefore be solved easily via a classical matrix-free conjugate gradient method. The final momentum is then given by (54a) and the scheme is completed by formally setting $\rho^{n+1,p}=\rho^{*,p}$, while the distorsion field and the thermal impulse are updated via the following compatible structure-preserving discretization that preserves curl-free fields exactly at the discrete level in the absence of source terms, by virtue of the use of compatible discrete nabla operators, see also [32, 35]:
@@
which is a compatible discretization of eqn. (1c), see also [32]. In order to enforce the compatibility with the determinant constraint we may uniformly rescale all components of $A_{ik}^{n+1,c}$ so that $|A_{ik}^{n+1,c}|=\rho^{n+1,c}/\rho_{0}$ holds. For alternative compatible discretizations, see [75, 74].
$J_{k}^{n+1,c}=J_{k}^{n,c}-\Delta t\,\partial_{k}^{cp}\left(v_{m}^{n+1,p}J_{m}^{n,p}+T^{**,p}\right)-\frac{\Delta t}{4}\sum\limits_{p\in\mathcal{P}_{c}}v_{m}^{n+1,p}\left(\partial_{m}^{pa}J_{k}^{n,a}-\partial_{k}^{pa}J_{m}^{n,a}\right)+\frac{\Delta t}{\theta_{2}}\beta_{ik}^{n+1,c},$ (57)
which is a compatible discretization of eqn. (1d), see also [32]. The discrete total energy is finally updated as
@@ 3.5 Summary of the algorithm
In the following we summarize the complete algorithm of the proposed structure-preserving four-split scheme.
@@ 3.5.2 Temperature subsystem
$\mathcal{E}^{**,c}=\mathcal{E}^{*,c}-\Delta t\,\partial_{k}^{cp}q_{k}^{**,p}.$ (63)
@@ 3.5.5 Compatible update of the distortion field, of the thermal impulse and of the total energy
In the absence of source terms the distortion field and the thermal impulse must remain curl-free for all times if they were initially curl-free. This is achieved via a compatible discretization of the original equations for $A_{ik}$ and $J_{k}$, i.e. eqns. (1c) and (1d), see also [32]:
@@
$J_{k}^{n+1,c}=J_{k}^{n,c}-\Delta t\,\partial_{k}^{cp}\left(v_{m}^{n+1,p}J_{m}^{n,p}+T^{**,p}\right)-\frac{\Delta t}{4}\sum\limits_{p\in\mathcal{P}_{c}}v_{m}^{n+1,p}\left(\partial_{m}^{pa}J_{k}^{n,a}-\partial_{k}^{pa}J_{m}^{n,a}\right)+\frac{\Delta t}{\theta_{2}}\beta_{ik}^{n+1,c}.$ (70)
@@
The proposed scheme is a finite volume method and is therefore consistent with the integral form of the conservation laws. Indeed, it conserves mass, momentum and total energy exactly at the discrete level. Unfortunately, we were not yet able to prove a discrete entropy inequality for this scheme, which is clearly a shortcoming. Further research in this direction will be the subject of future work.
@@ 4.4 Lid-driven cavity
Compared to a classical explicit FV scheme our timestep is 10,000 times larger, and compared to the structure-preserving semi-implicit scheme introduced in [32], in which shear and heat waves were still treated explicitly, the timestep of our new scheme is 1000 times larger.
@@ 4.5 Solid rotor
The aim of this test case is to numerically verify the discrete curl-free property of our new 4-split semi-implicit finite volume scheme. For the setup of the test case we follow [32], solving (1) with $\tau_{1}=10^{20}$, but here we set $\tau_{2}=10^{-14}$, i.e. the strain relaxation source term is absent, while heat conduction is considered in the stiff Fourier limit. The computational domain is given by $\Omega=[-1,+1]^{2}$. The initial condition is $\rho=1$, $p=10^{5}$, $\mathbf{A}=\mathbf{I}$ and $\mathbf{J}=0$ and the initial velocity field is u=y/Ru=-y/R, $v=+x/R$ and $w=0$ within the circular region $r\leq R$, where $r=\left\|\mathbf{x}\right\|_{2}$ and $R=0.2$, while $\mathbf{v}=0$ for $r>R$. The remaining parameters are $\gamma=1.4$, $c_{v}=1004/\gamma$, $c_{s}=1.0$ and $c_{h}=100$. With this setting we are in the low Mach regime for the acoustic Mach number $M_{a}$ and for the heat Mach number $M_{h}$. The test case is solved with the new structure-preserving semi-implicit 4-split finite volume scheme until a final time of $t=0.3$. The computational mesh uses $512\times 512$ cells. In order to obtain a numerical reference solution, the same problem is solved on the same grid with the second order version of the 2-split structure-preserving scheme introduced in [32], but in the absence of heat conduction, i.e. setting $c_{h}=0$, since the 2-split method presented in [32] is not able to deal with low heat Mach numbers $M_{h}$. The results obtained with the new 4-split scheme are depicted in Fig. 9, where the contour colors of the velocity component $u$ and the distortion field components $A_{11}$ and $A_{12}$ are shown. A comparison with the reference solution is provided in the 1D cut shown in the left panel of Fig. 10, while in the right panel of Fig. 10 we show the the time series of the curl errors of $\mathbf{A}$ and $\mathbf{J}$ in the $L^{\infty}$ norm. One can observe a good agreement with the reference solution and that the new structure-preserving semi-implicit 4-split scheme presented in this paper produces curl errors that remain at the order of machine precision, as expected. A standard MUSCL-Hancock scheme produces instead larger curl errors that are of the order of the mesh spacing, see [32] for comparison. In Figure 11 we report the temporal evolution of mass, $x$-momentum, entropy and of the kinetic energy as well as of the energy stored in the deformation of the solid. For this smooth problem, entropy is preserved, as expected.
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 11: Solid rotor problem. Time evolution of mass, $x$-momentum, kinetic and deformation energy and entropy.
@@ 4.6 Circular explosion problem
We first solve the governing PDE system in the quasi inviscid fluid limit of the model, i.e. $\tau_{1}\ll 1$ and $\tau_{2}\ll 1$. For the inner state we set the density and the pressure to $\rho_{in}=1$ and $p_{in}=1$, while in the outer state we impose $\rho_{out}=0.125$ and $p_{out}=0.1$. The initial velocity is globally set to $\mathbf{v}=0$, the initial thermal impulse vector is equal to $\mathbf{J}=0$ and the distortion field is initialized as $\mathbf{A}=\mathbf{I}$. The other parameters are chosen as $\gamma=1.4$, $c_{v}=2.5$, $c_{s}=1$, $c_{h}=1$, $\rho_{0}=1$, $\tau_{1}=10^{-8}$ and $\tau_{2}=10^{-10}$. The mesh is composed by $500\times 500$ control volumes and the final time is set to $t=0.2$. Note that with the chosen parameters, all Mach numbers are of the order of unity. The reference solution is obtained by solving a 1D radial Euler system with source terms using a classical second-order MUSCL-Hancock scheme on a very fine mesh, as described in detail in [80, 32]. In the left panel of Fig. 12 the 3D density contours are shown at the final time. In Fig. 13 a comparison of the reference solution with the numerical results obtained at the aid of the new semi-implicit 4-split scheme is provided. One observes overall a good agreement. In Fig. 14 we present the time evolution of mass, $x$-momentum, total energy and entropy integrated over the domain. As expected, mass momentum and total energy are perfectly conserved. Our numerical results show that physical entropy is monotonically increasing in time, as it should, although we were not able to prove a discrete entropy inequality for our scheme.
@@
We now solve the governing PDE system in the solid limit, i.e. $\tau_{1}\to\infty$ and $\tau_{2}\to\infty$, with the same initial conditions and parameters as described above for EP1, apart from the relaxation times, which are set to $\tau_{1}=\tau_{2}=10^{20}$. The mesh is again composed of $500\times 500$ cells and the final time is now set to $t=0.15$. We solve the problem twice, once with the new 4-split SPSIFV scheme and another time with the thermodynamically compatible HTC scheme presented in [68, 82], in order to obtain a numerical reference solution. A 3D density contour plot at the final time is depicted in the right panel of Fig. 12. The comparison of the computational results between the new 4-split SIFV scheme and the numerical reference solution is shown in via 1D cuts along the $x$-axis in Fig. 15. One can again note a good agreement between the two solutions. In Fig. 16 we show the time evolution of mass, $x$-momentum, total energy and entropy integrated over the domain. Again, mass momentum and total energy are perfectly conserved, while physical entropy is monotonically increasing in time, as it should be. However, we have to stress again that so far we were not able to prove an entropy inequality for the scheme presented in this paper.
@@
Figure 14: 2D explosion problem EP1. Time evolution of mass, $x$-momentum, total energy and entropy.
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 16: 2D explosion problem EP2. Time evolution of mass, $x$-momentum, total energy and physical entropy.
@@ 5 Conclusions
Further research will be devoted to the adoption of unstructured meshes with the aim of tackling more complex geometries, while preserving additional structure of the system exactly, see e.g. [75, 73, 74], as well as the incorporation of electromagnetic fields, following the general lines of [83]. To improve the accuracy in space, high order extension is also foreseen following the ideas recently forwarded in [36], while higher order in time may be achieved according to the lines presented in [84]. Last but not least, in future work we will try to find a compatible semi-implicit discretization that also provably satisfies a cell entropy inequality.